# Multi-Robot Connected Fermat Spiral Coverage

### v1

@@ Abstract
We introduce the Multi-Robot Connected Fermat Spiral (MCFS), a novel algorithmic framework for Multi-Robot Coverage Path Planning (MCPP) that adapts Connected Fermat Spiral (CFS) from the computer graphics community to multi-robot coordination for the first time. MCFS uniquely enables the orchestration of multiple robots to generate coverage paths that contour around arbitrarily shaped obstacles, a feature that is notably lacking in traditional methods. Our framework not only enhances area coverage and optimizes task performance, particularly in terms of makespan, for workspaces rich in irregular obstacles but also addresses the challenges of path continuity and curvature critical for non-holonomic robots by generating smooth paths without decomposing the workspace. MCFS solves MCPP by constructing a graph of isolines and transforming MCPP into a combinatorial optimization problem, aiming to minimize the makespan while covering all vertices. Our contributions include developing a unified CFS version for scalable and adaptable MCPP, extending it to MCPP with novel optimization techniques for cost reduction and path continuity and smoothness, and demonstrating through extensive experiments that MCFS outperforms existing MCPP methods in makespan, path curvature, coverage ratio, and overlapping ratio. Our research marks a significant step in MCPP, showcasing the fusion of computer graphics and automated planning principles to advance the capabilities of multi-robot systems in complex environments. Our code is available at https://github.com/reso1/MCFS.
@@ 2 Related Work
Grid-Based Methods: Grid-based coverage methods abstract workspaces into square grids (Hazon and Kaminka 2005; Kapoutsis, Chatzichristofis, and Kosmatopoulos 2017; Tang, Sun, and Zhang 2021), allowing for the application of various graph algorithms. One prominent method, Spanning Tree Coverage (STC) (Gabriely and Rimon 2001), constructs a minimum spanning tree and then generates circumnavigating paths on the tree to cover the workspace. STC-based MCPP methods (Hazon and Kaminka 2005; Tang and Ma 2023) work by finding a set of trees that jointly visit all vertices and assigning each robot a path that circumnavigates a tree. While convenient, the complexity of optimally solving grid-based MCPP grows exponentially in the workspace size and the number of robots.
@@ 3 Connected Fermat Spiral (CFS)
3.1 Constructing Isolines and the Isograph
@@
where $d(\mathbf{p},I)$ denotes the distance between point $\mathbf{p}$ and isoline $I$. Unlike the original CFS which directly constructs an undirected edge $(u,v)$ if $O_{u\rightarrow v}$ is nonempty, we also consider $O_{v\rightarrow u}$ for edge construction. This consideration provides flexibility in traversing the isograph in any order and from any root isovertex in the CFS context. It also avoids adding edges $(u,v)$ where the respective isolines $I_{u}$ and $I_{v}$ are separated by multiple isolines, as such pairs may be unsuitable for stitching in the CPP context (see the case study in Sec. 3.4). Therefore, we define a set $O_{u,v}$ of stitching tuples for any $u,v\in V$ in adjacent layers as:
@@
3.2 Unifying the CFS algorithm
We detail our unified version of CFS in Alg. 1, which takes as input an isograph $G$ and an entry point $\mathbf{p}_{0}$. The algorithm starts by identifying the isovertex $r$ containing $\mathbf{p}_{0}$ [Line 1] as the root for a depth-first search (DFS) traversal of $G$. It then initializes the CFS path $\pi$ to be constructed and the set $U$ to record points already used to stitch the isolines [Line 1]. The main loop then iterates over the DFS edges [Line 1] and stitches the corresponding pair of isolines for each edge [Lines 1-1]. Specifically, a stitching tuple $(\mathbf{p},\mathbf{q})$ is selected via any selector [Line 1]. For any isovertex $v\in V$, we use $I_{v}(\mathbf{p})$ to denote the counterclockwise path along isoline $I_{v}$ starting at $\mathbf{p}$ and ending at $\mathcal{B}_{v}(\mathbf{p})$. This path segment is then stitched into $\pi$ using the selected stitching tuple [Line 1]. The set $U$ is updated to include these newly selected stitching tuples [Line 1]. Following the iterations over all DFS edges, the final path $\pi$ is constructed to stitch together all isolines and completely cover the given polygon.
@@ 3.2 Unifying the CFS algorithm
1 $r\leftarrow$ the isovertex of $G$ containing $\mathbf{p}_{0}$ $\pi\leftarrow I_{r}(\mathbf{p}_{0}),\;U\leftarrow\emptyset$ for $(u,v)\in$ DFS traversal edges of $G$ from $r$ do
2    remove any $(\mathbf{p},\mathbf{q})$ from $O_{u,v}$ where $\mathbf{p}\in U$ or $\mathbf{q}\in U$ $(\mathbf{p},\mathbf{q})\leftarrow f\left(O_{u,v}\right)$triangleright by any stitching tuple selector $f$ stitch $I_{v}(\mathbf{p})$ into $\pi$ by stitching $\mathbf{p}$ to $\mathbf{q}$ and $\mathcal{B}(\mathbf{p})$ to $\mathcal{B}(\mathbf{q})$ $U\leftarrow U\cup\{\mathbf{p},\mathbf{q}\}$
return $\pi$
@@
3.3 Stitching Tuple Selector
@@
3.4 Case Study: Unified vs Original CFS
@@ 4 Multi-Robot CFS Coverage
In this section, we present our MCFS framework for solving MCPP. MCFS computes multiple trees from an input isograph, each corresponding to a different robot, and then applies CFS on each tree to compute individual coverage paths. In Sec. 4.1, we detail the CFS-based formulation of MCPP and introduce its reduction to Min-Max Rooted Tree Cover (MMRTC) (Even et al. 2004; Tang and Ma 2023). We then present two optimization techniques, isograph augmentation in Sec. 4.2 and MMRTC solution refinement in Sec. 4.3, aiming to further enhance the MCPP solution.
4.1 Problem Formulation
@@
When using CFS to generate each coverage path in $\Pi$, the path length is linear in $|\pi|$ and therefore the cost of any path $\pi$ can be evaluated as $c(\pi)=|\pi|$, since each isoline in CFS contains equidistant points (as detailed in Sec.3.1). For an isograph $G=(V,E)$, each $v\in V$ is assigned a weight $w_{v}=|I_{v}|$, representing the number or points in isoline $I_{v}$. Consequently, the cost of any tree $T\subseteq G$ is $c(T)=\sum_{v\in V(T)}w_{v}$. The MMRTC problem parallels MCPP in its aim of finding a makespan-minimizing set of rooted trees, where each graph vertex is covered by at least one tree. Given a graph $G=(V,E)$ and a set $R=\{r_{i}\}_{i\in I}\subseteq V$ of root isovertices for robots, the objective of MMRTC is defined as:
@@
where each $T_{i}\in\mathcal{T}$ is a tree rooted at $r_{i}$ and $c(T_{i})$ is its tree cost. Let $V(T)$ and $E(T)$ denote the vertex set and edge set of any tree $T$, respectively. The solution set $\mathcal{T}$ must satisfy $v\in\bigcup_{i\in I}V(T_{i})$ to ensure the coverage of all $v\in V$. Since CFS stitches each isoline $I_{v}$ of $v\in V(T_{i})$ to construct the coverage path $\pi_{i}\in\Pi$, we have $c(\pi_{i})=|\pi_{i}|=\sum_{v\in V(T_{i})}|I_{v}|={c(T_{i})}$. Hence, for any isograph $G$ and set $R$ of root isovertices for robots, the heuristic values in Eqn. (4) and Eqn. (5) are identical under CFS, effectively reducing MCPP to MMRTC.
We employ the Mixed Integer Programming (MIP) model proposed in (Tang and Ma 2023) to solve MMRTC optimally. The optimal set of trees obtained is then used to produce coverage paths by applying our unified CFS (Alg. 1) on each tree. Fig. 4-(a) and (b) illustrate a 2-tree MMRTC instance and its corresponding solution. The details of the MIP model are shown in Appendix B.
4.2 Optimization: Isograph Augmentation
@@
where $d_{G}(\cdot,\cdot)$ denotes the graph distance between any two isovertices in $G$, and $\delta$ is a hyperparameter that sets the augmentation level. For the edges in $E^{\#}$, stitching tuples are constructed differently from those edges in the original isograph edge set $E$. Without loss of generality, we consider an edge $(v_{1},v_{k+1})\in E^{\#}$ and its shortest path $(v_{1},v_{2},...,v_{k+1})$ in the original $G$ (i.e., each segment $(v_{i},v_{i+1})$ is part of $E$ and $k$ is the graph distance between $v_{1}$ and $v_{k+1}$). The set $O_{v_{1},v_{k+1}}$ comprises all pairs of $\mathbf{p}_{1}$ on the isoline of $v_{1}$ and $\mathbf{p}_{k+1}$ on the isoline of $v_{k+1}$ that can be feasibly connected, forming valid stitching tuples $(\mathbf{p}_{1},\mathbf{p}_{k+1})$. Such points are connectable iff they form a sequence of consecutive stitching tuples $(\mathbf{p}_{1},\mathbf{p}_{2})\in O_{v_{1},v_{2}},\ldots,(\mathbf{p}_{k},\mathbf{p}_{k+1})\in O_{v_{k},v_{k+1}}$, which ensures that the straight-line segment between the pair does not intersect more than $k-1$ or any obstacles within the workspace. Fig. 3 demonstrates an example of the adding procedure of an augmented edge $(v_{1},v_{3})$ with three valid stitching tuples in $O_{v_{1},v_{3}}$ and how $p_{1}$ and $p_{3}$ can be connected via $p_{2}$. Given that the distance between adjacent isolines is set as $l$ previously, we assign a weight $w_{e}=l\times k$ to each $e=(u,v)\in E^{\#}$ with a layer difference of $k$ (i.e., $|L_{u}-L_{v}|=k$), which approximates the additional path cost incurred by any tree containing $e$. The cost of any tree $T$ is thus updated to $c(T)=\sum_{v\in V(T)}w_{v}+\sum_{e\in E(T)}w_{e}$ in the MMRTC solving. Once the augmented edge set $E^{\#}$ and the corresponding stitching tuple sets $O$ are constructed, the original isograph $G$ is updated by setting $E=E\cup E^{\#}$, and the same MMRTC model is solved on the augmented $G$.
@@
1 optimized solution $\mathcal{T}^{*}\leftarrow\mathcal{T}$, set of used isovertices $U\leftarrow\emptyset$ $M\leftarrow$ set of all repeatedly visited isovertices in $\mathcal{T}$ call AddImprovingRepetition($\mathcal{T},M,U$) if $M=\emptyset$ max-heapify $M$ ordered by the number of occurrences while $M\neq\emptyset$ do
2    $u\leftarrow M.pop()$ $\mathcal{T}_{u}\leftarrow$ set of all trees containing $u$ in current solution $\mathcal{T}$ for $(u,v)\in\{(u,v)\in E\,|\,v\notin U\}$ do
3        $h,\mathcal{T}_{u}\leftarrow\textsc{PairwiseIsoverticesSplitting}(\mathcal{T}_{u},u,v)$ set $h^{*}$ to $h$ and $\mathcal{T}^{*}_{u}$ to $\mathcal{T}_{u}$ if $h<h^{*}$
4    use $\mathcal{T}_{u}^{*}$ to update $\mathcal{T}$, $U\leftarrow U\cup\{u,v\}$, $M\leftarrow M/\{v\}$ set $\mathcal{T}^{*}$ to $\mathcal{T}$ if its evaluated makespan is smaller call AddImprovingRepetition($\mathcal{T},M,U$) if $M=\emptyset$
5return $\mathcal{T}^{*}$ Function AddImprovingRepetition($\mathcal{T},M,U$):
6    $P\leftarrow$ set of leaf isovertices (i.e., with a degree of $1$) $u\notin U$ in the highest-cost tree in $\mathcal{T}$ that are not from PIS splitting $T,u\leftarrow$ lowest-cost $T\in\mathcal{T}$ and any $u\in P$ such that $u$ is not in $T$ but a neighbor of some $v\notin U$ in $T$ add $u$ and edge $(u,v)$ to $T$, $M\leftarrow M\cup\{u\}$
@@ 7 Function PairwiseIsoverticesSplitting($\mathcal{T}_{u},u,v$):
8    $h^{*}\leftarrow+\infty,\mathcal{T}^{*}_{u}\leftarrow\mathcal{T}_{u}$ for $\mathbf{o}=(o_{1},\ldots,o_{|\mathcal{T}_{u}|})\in O_{u,v}^{|\mathcal{T}_{u}|}$ do
9        $\mathcal{T}_{u}^{\prime}\leftarrow$ a copy of $\mathcal{T}_{u}$ split $u,v$ into $|\mathcal{T}_{u}|$ new isovertices by stitching $I_{u},I_{v}$ via $\mathbf{o}$, each assigned to a $T\in\mathcal{T}_{u}^{\prime}$triangleright see Fig. 4-(b)(c)(d) for $T\in\mathcal{T}_{u}$ do
10
add the assigned new isovertex $z$ to $T$ and remove $u$ from $T$
replace each edge $(u,\cdot)$ with edge $(z,\cdot)$ in $T$ if $v\in V(T)$ then
11                remove $v$ from $T$ and replace each $(v,\cdot)$ with $(z,\cdot)$ in $T$

### v3

@@ Abstract
We introduce the Multi-Robot Connected Fermat Spiral (MCFS), a novel algorithmic framework for Multi-Robot Coverage Path Planning (MCPP) that adapts Connected Fermat Spiral (CFS) from the computer graphics community to multi-robot coordination for the first time. MCFS uniquely enables the orchestration of multiple robots to generate coverage paths that contour around arbitrarily shaped obstacles, a feature that is notably lacking in traditional methods. Our framework not only enhances area coverage and optimizes task performance, particularly in terms of makespan, for workspaces rich in irregular obstacles but also addresses the challenges of path continuity and curvature critical for non-holonomic robots by generating smooth paths without decomposing the workspace. MCFS solves MCPP by constructing a graph of isolines and transforming MCPP into a combinatorial optimization problem, aiming to minimize the makespan while covering all vertices. Our contributions include developing a unified CFS version for scalable and adaptable MCPP, extending it to MCPP with novel optimization techniques for cost reduction and path continuity and smoothness, and demonstrating through extensive experiments that MCFS outperforms existing MCPP methods in makespan, path curvature, coverage ratio, and overlapping ratio. Our research marks a significant step in MCPP, showcasing the fusion of computer graphics and automated planning principles to advance the capabilities of multi-robot systems in complex environments. Our code is publicly available at https://github.com/reso1/MCFS.
@@ 2 Related Work
Grid-Based Methods: Grid-based coverage methods abstract workspaces into square grids (Hazon and Kaminka 2005; Kapoutsis, Chatzichristofis, and Kosmatopoulos 2017; Tang, Sun, and Zhang 2021), allowing for the application of various graph algorithms. One prominent method, Spanning Tree Coverage (STC) (Gabriely and Rimon 2001), constructs a minimum spanning tree and then generates circumnavigating paths on the tree to cover the workspace. STC-based MCPP methods (Hazon and Kaminka 2005; Tang and Ma 2023, 2024) work by finding a set of trees that jointly visit all vertices and assigning each robot a path that circumnavigates a tree. While convenient, the complexity of optimally solving grid-based MCPP grows exponentially in the workspace size and the number of robots.
@@ 3 Connected Fermat Spiral (CFS)
Constructing Isolines and the Isograph
@@
where $d(\mathbf{p},I)$ denotes the distance between point $\mathbf{p}$ and isoline $I$. Unlike the original CFS which directly constructs an undirected edge $(u,v)$ if $O_{u\rightarrow v}$ is nonempty, we also consider $O_{v\rightarrow u}$ for edge construction. This consideration provides flexibility in traversing the isograph in any order and from any root isovertex in the CFS context. It also avoids adding edges $(u,v)$ where the respective isolines $I_{u}$ and $I_{v}$ are separated by multiple isolines, as such pairs may be unsuitable for stitching in the CPP context (see Fig. 2 for the case study). Therefore, we define a set $O_{u,v}$ of stitching tuples for any $u,v\in V$ in adjacent layers as:
@@
1 $r\leftarrow$ the isovertex of $G$ containing $\mathbf{p}_{0}$
2 $\pi\leftarrow I_{r}(\mathbf{p}_{0}),\;U\leftarrow\emptyset$
3 for $(u,v)\in$ DFS traversal edges of $G$ from $r$ do
4    remove any $(\mathbf{p},\mathbf{q})$ from $O_{u,v}$ where $\mathbf{p}\in U$ or $\mathbf{q}\in U$
5    $(\mathbf{p},\mathbf{q})\leftarrow f\left(O_{u,v}\right)$triangleright by any stitching tuple selector $f$
6    stitch $I_{v}(\mathbf{p})$ into $\pi$ by stitching $\mathbf{p}$ to $\mathbf{q}$ and $\mathcal{B}(\mathbf{p})$ to $\mathcal{B}(\mathbf{q})$
7    $U\leftarrow U\cup\{\mathbf{p},\mathbf{q}\}$
8
9return $\pi$
@@ 7    $U\leftarrow U\cup\{\mathbf{p},\mathbf{q}\}$
Unifying the CFS Algorithm
We detail our unified version of CFS in Alg. 1, which takes as input an isograph $G$ and an entry point $\mathbf{p}_{0}$. The algorithm starts by identifying the isovertex $r$ containing $\mathbf{p}_{0}$ [Line 1] as the root for a depth-first search (DFS) traversal of $G$. It then initializes the CFS path $\pi$ to be constructed and the set $U$ to record the points already used to stitch the isolines [Line 1]. The main loop then iterates over the DFS edges [Line 1] and stitches the corresponding pair of isolines for each edge [Lines 1-1]. Specifically, a stitching tuple $(\mathbf{p},\mathbf{q})$ is selected via any selector [Line 1]. For any isovertex $v\in V$, we use $I_{v}(\mathbf{p})$ to denote the counterclockwise path along isoline $I_{v}$ starting at $\mathbf{p}$ and ending at $\mathcal{B}_{v}(\mathbf{p})$. This path segment is then stitched into $\pi$ using the selected stitching tuple [Line 1]. The set $U$ is updated to include these newly selected stitching tuples [Line 1]. Following the iterations over all DFS edges, the final path $\pi$ is constructed to stitch together all isolines and completely cover the given polygon.
@@
Stitching Tuple Selector
@@
Case Study: Unified vs Original CFS
@@ 4 Multi-Robot CFS Coverage
In this section, we present our MCFS framework for solving MCPP. MCFS computes multiple trees from an input isograph, each corresponding to a different robot, and then applies CFS on each tree to compute individual coverage paths. First, we detail the CFS-based formulation of MCPP and introduce its reduction to Min-Max Rooted Tree Cover (MMRTC) (Even et al. 2004; Tang and Ma 2023). We then present two optimization techniques, isograph augmentation and MMRTC solution refinement, aiming to further enhance the MCPP solution.
Problem Formulation
@@
When using CFS to generate each coverage path in $\Pi$, the path length is linear in $|\pi|$ and therefore the cost of any path $\pi$ can be evaluated as $c(\pi)=|\pi|$, since each isoline in CFS contains equidistant points (as detailed in the last section). For an isograph $G=(V,E)$, each $v\in V$ is assigned a weight $w_{v}=|I_{v}|$, representing the number or points in isoline $I_{v}$. Consequently, the cost of any tree $T\subseteq G$ is $c(T)=\sum_{v\in V(T)}w_{v}$. The MMRTC problem parallels MCPP in its aim of finding a makespan-minimizing set of rooted trees, where each graph vertex is covered by at least one tree. Given a graph $G=(V,E)$ and a set $R=\{r_{i}\}_{i\in I}\subseteq V$ of root isovertices for robots, the objective of MMRTC is defined as:
@@
where each $T_{i}\in\mathcal{T}$ is a tree rooted at $r_{i}$ and $c(T_{i})$ is its tree cost. Let $V(T)$ and $E(T)$ denote the vertex set and edge set of any tree $T$, respectively. The solution set $\mathcal{T}$ must satisfy $v\in\bigcup_{i\in I}V(T_{i})$ to ensure the coverage of all $v\in V$. Since the CFS stitches each isoline $I_{v}$ of $v\in V(T_{i})$ to construct the coverage path $\pi_{i}\in\Pi$, we have $c(\pi_{i})=|\pi_{i}|=\sum_{v\in V(T_{i})}|I_{v}|={c(T_{i})}$. Therefore, for any isograph $G$ and the set $R$ of root isovertices for robots, the objective values in Eqn. (4) and Eqn. (5) are identical under CFS, effectively reducing MCPP to MMRTC.
We employ the Mixed Integer Programming (MIP) model proposed in (Tang and Ma 2023) to solve MMRTC optimally. The optimal set of trees obtained is then used to produce coverage paths by applying our unified CFS (Alg. 1) on each tree. Figs. 4-(a) and (b) illustrate a 2-tree MMRTC instance and its corresponding solution. The details of the MIP model are shown in Appendix B.
Optimization: Isograph Augmentation
@@
where $d_{G}(\cdot,\cdot)$ denotes the graph distance between any two isovertices in $G$, and $\delta$ is a hyperparameter that sets the augmentation level. For the edges in $E^{\#}$, stitching tuples are constructed differently from those edges in the original isograph edge set $E$. Without loss of generality, we consider an edge $(v_{1},v_{k+1})\in E^{\#}$ and its shortest path $(v_{1},v_{2},...,v_{k+1})$ in the original $G$ (i.e., each segment $(v_{i},v_{i+1})$ is part of $E$ and $k$ is the graph distance between $v_{1}$ and $v_{k+1}$). The set $O_{v_{1},v_{k+1}}$ comprises all pairs of $\mathbf{p}_{1}$ on the isoline of $v_{1}$ and $\mathbf{p}_{k+1}$ on the isoline of $v_{k+1}$ that can be feasibly connected, forming valid stitching tuples $(\mathbf{p}_{1},\mathbf{p}_{k+1})$. Such points are connectable iff they form a sequence of consecutive stitching tuples $(\mathbf{p}_{1},\mathbf{p}_{2})\in O_{v_{1},v_{2}},\ldots,(\mathbf{p}_{k},\mathbf{p}_{k+1})\in O_{v_{k},v_{k+1}}$, which ensures that the straight-line segment between the pair does not intersect more than $k-1$ isoline(s) or any obstacles within the workspace. Fig. 3 demonstrates an example of the adding procedure of an augmented edge $(v_{1},v_{3})$ with three valid stitching tuples in $O_{v_{1},v_{3}}$ and how $p_{1}$ and $p_{3}$ can be connected via $p_{2}$. Given that the distance between adjacent isolines is set as $l$ previously, we assign a weight $w_{e}=l\times k$ to each $e=(u,v)\in E^{\#}$ with a layer difference of $k$ (i.e., $|L_{u}-L_{v}|=k$), which approximates the additional path cost incurred by any tree containing $e$. The cost of any tree $T$ is thus updated to $c(T)=\sum_{v\in V(T)}w_{v}+\sum_{e\in E(T)}w_{e}$ in the MMRTC solving. Once the augmented edge set $E^{\#}$ and the corresponding stitching tuple sets $O$ are constructed, the original isograph $G$ is updated by setting $E=E\cup E^{\#}$, and the same MMRTC model is solved on the augmented $G$.
@@
1 optimized solution $\mathcal{T}^{*}\leftarrow\mathcal{T}$, set of used isovertices $U\leftarrow\emptyset$
2 $M\leftarrow$ set of all repeatedly visited isovertices in $\mathcal{T}$
3 call AddImprovingRepetition($\mathcal{T},M,U$) if $M=\emptyset$
4 max-heapify $M$ ordered by the number of occurrences
5 while $M\neq\emptyset$ do
6    $u\leftarrow M.pop()$
7    $\mathcal{T}_{u}\leftarrow$ set of all trees containing $u$ in current solution $\mathcal{T}$
8    for $(u,v)\in\{(u,v)\in E\,|\,v\notin U\}$ do
9        $h,\mathcal{T}_{u}\leftarrow\textsc{PairwiseIsoverticesSplitting}(\mathcal{T}_{u},u,v)$
10        set $h^{*}$ to $h$ and $\mathcal{T}^{*}_{u}$ to $\mathcal{T}_{u}$ if $h<h^{*}$
11
12   use $\mathcal{T}_{u}^{*}$ to update $\mathcal{T}$, $U\leftarrow U\cup\{u,v\}$, $M\leftarrow M/\{v\}$
13    set $\mathcal{T}^{*}$ to $\mathcal{T}$ if its evaluated makespan is smaller
14    call AddImprovingRepetition($\mathcal{T},M,U$) if $M=\emptyset$
15
16return $\mathcal{T}^{*}$
17 Function AddImprovingRepetition($\mathcal{T},M,U$):
18    $P\leftarrow$ set of leaf isovertices (i.e., with a degree of $1$) $u\notin U$ in the highest-cost tree in $\mathcal{T}$ that are not from PIS splitting
19    $T,u\leftarrow$ lowest-cost $T\in\mathcal{T}$ and any $u\in P$ such that $u$ is not in $T$ but a neighbor of some $v\notin U$ in $T$
20    add $u$ and edge $(u,v)$ to $T$, $M\leftarrow M\cup\{u\}$
21
@@ 22 Function PairwiseIsoverticesSplitting($\mathcal{T}_{u},u,v$):
23    $h^{*}\leftarrow+\infty,\mathcal{T}^{*}_{u}\leftarrow\mathcal{T}_{u}$
24    for $\mathbf{o}=(o_{1},\ldots,o_{|\mathcal{T}_{u}|})\in O_{u,v}^{|\mathcal{T}_{u}|}$ do
25        $\mathcal{T}_{u}^{\prime}\leftarrow$ a copy of $\mathcal{T}_{u}$
26        split $u,v$ into $|\mathcal{T}_{u}|$ new isovertices $z$’s by stitching $I_{u},I_{v}$ via $\mathbf{o}$, each assigned to a $T\in\mathcal{T}_{u}^{\prime}$triangleright see Figs. 4-(b)(c)(d)
27        for $T\in\mathcal{T}_{u}$ do
28            if $v\in T$ then
29                replace each edge $(u,\cdot)$ or $(v,\cdot)$ (except edge $(u,v)$) with $(z,\cdot)$ in $T$ and remove $u,v$ from $T$
30
31           else
32               replace each $(u,\cdot)$ with $(z,\cdot)$ in $T$ and remove $u$ from $T$